OpenStudy (anonymous):

completely factor the trinomial or indicate if its not factorable: -3cq^3-3cq^2-15cq

OpenStudy (anonymous):

-3cq(q^2+q+5) for q^2+q+5 d=1-20=-19<0 it is not factorable